OpenStudy (anonymous):

\[A\] is a subset of \[\mathbb{R} \quad \Large{A:=\{\frac{1}{n^{2}+4} \quad | \quad n \in \mathbb{N}_{+}\}}\] determine wheter set is bounded from above or below and deterimine if there is sup or inf exists

OpenStudy (anonymous):

the set is bounded below

OpenStudy (anonymous):

can you prove it pls ?

OpenStudy (anonymous):

txh loser but i need more detailed description

OpenStudy (anonymous):

I need a prove like by attachment

OpenStudy (anonymous):

ok that sounds better, thank you loser ;)

OpenStudy (anonymous):

its weird when you say loser

OpenStudy (loser66):

@jonask I don't know, I was taught that, if it's weird to you, so, I take it off .

OpenStudy (anonymous):

hey Loser whats problem i need the solution pls add back ;) i think there is some missunderstanding between you and jonask

OpenStudy (loser66):

@tunahan let me tag other for help. I am not good at explanation. @cwrw238

OpenStudy (loser66):

please!!

OpenStudy (anonymous):

i was happy with your explanation :)

OpenStudy (loser66):

let wait for hartn. he is good at explanation. I knew it

OpenStudy (anonymous):

ok good :)

OpenStudy (loser66):

As diverge test states that lim A_n when n goes to infinitive is not 0, then A_n diverges. therefore, I take lim A_n = lim (1/n^2 +4) =0 when n goes to infinitive. That includes that A_n converges to 0 , In other words, 0 is greatest low bound of A_n I don't delete it.

OpenStudy (anonymous):

thx

OpenStudy (loser66):

@hartnn think of this, if I have A_n is set of numbers such that A_n = 1/n^2 +4. if I have n=1 , A_1 = 1/5 n=2, A_2 = 1/8 n=3 A_3 = 1/13 so on and so on .... so that my subset A goes to where?? A ={ 1/5,1/8/,1/13.......} it goes closer and closer to 0. Although it never get 0 , but 0 is the point it will come up. think about the function down, down down...... but never negative. therefore, line 0 is the greatest low bound.